# Less is More: Skim Transformer for Light Field Image Super-resolution

### v2

@@ IV-E3 Ablation study of Skimmed SAI Sets
Column (g) experiments with a minimal variant using only a single central SAI (yellow). In this setting, the angular scope of the attention score matrix is severely restricted, resulting in a substantial performance drop of 0.53 dB compared to the baseline. This highlights the necessity of the disparity modeling capabilities offered by DSA.
@@
While the previous cases reduce either the number of DSA branches or the number of SAIs, each causing performance degradation, we find that merely increasing the number of SAIs does not necessarily lead to better results. Column (h) augments the baseline with a third DSA branch comprising central SAIs from each edge (green). Although this triple-branch variant introduces additional parameters and computational cost, it does not yield any noticeable performance improvement. This outcome suggests that the baseline configuration already provides sufficient disparity range coverage.
In a more extreme case, Column (i) consolidates all SAIs into a single DSA branch. This effectively reintroduces disparity entanglement, as the single branch must process all the SAI at once. Despite a 61% increase in model parameters, this configuration results in a 0.08 dB decrease in performance. This further confirms the importance of the information structure constructed by DSA branches over merely the amount of information involved.
Collectively, these results underscore the flexibility and effectiveness of the skimmed SAI design in DSA branches. More importantly, they highlight the need for strategic SAI configuration to achieve comprehensive disparity coverage with minimal computational overhead, achieving “less is more”.
V Conclusion
@@
In this paper, we presented SkimLFSR, a novel and efficient LFSR network that overcomes the challenge of disparity entanglement in LF image processing. The experimental results validated SkimLFSR’s “less is more” philosophy, demonstrating that SkimLFSR outperformed the state-of-the-art methods significantly by utilizing less but more pertinent information with reduced memory and computational complexity. Several in-depth analyses were conducted to reveal the underlying disparity-aware mechanisms of SkimLFSR.
Despite its success, our work has certain limitations. First, the disparity ranges for the Skim Transformers must be manually pre-specified. Although this is not a complex problem, it essentially introduces a prior into the network, which could potentially be made heuristic or learned in network optimization autonomously. Second, the LFSR datasets used in this study have limited angular resolution ($\times 5$ 5), which does not fully showcase SkimLFSR’s potential. Nevertheless, we believe that SkimLFSR could serve as a new paradigm for not only LFSR but also other LF-related tasks, allowing it to handle more complex angular resolutions.

### v3

@@ 2. Ablation Study on Skimmed SAI Sets (Trained on $7\times 7$ SAIs)
Variant (g) experiments with a minimal variant using only a single central SAI (yellow). In this setting, the angular scope of the attention score matrix is severely restricted, resulting in a substantial performance drop of 0.53 dB compared to the baseline. This highlights the necessity of the multi-branch disparity modeling.
@@
While the previous cases reduce either the number of DSA branches or the number of SAIs, each causing performance degradation, we find that merely increasing the number of SAIs does not necessarily lead to better results. Variant (h) augments the baseline with a third DSA branch comprising central SAIs from each edge (green). Although this triple-branch variant introduces additional parameters and computational cost, it does not yield any noticeable performance improvement. This outcome suggests that the baseline configuration already provides sufficient disparity range coverage.
In a more extreme case, Variant (i) consolidates all SAIs into a single DSA branch. This effectively reintroduces disparity entanglement back to the model, as the single branch must process all the SAI at once. Despite a slight performance drop of 0.08 dB and a similar FLOPs and inference time, this all-in-one configuration results in a 61% increase in model parameters. This increase in parameters not only introduces additional computational burden but also raises the risk of overfitting and compromises the model’s scalability. To further investigate this effect, we compare the baseline Variant (a) with the all-in-one Variant (i) by varying the number of correlation blocks $N_{CB}$ from 4 to 20 in increments of two. The results, shown in Fig. 9, reveal that Variant (i) not only underperforms the baseline across all $N_{CB}$ but also saturates early at $N_{CB}=16$ with only 32.76 dB PSNR, whereas the baseline continues to improve and reaches 33.06 dB at $N_{CB}=20$.
Collectively, these findings confirm that the information structure introduced by the DSA branches is far more important than merely increasing the amount of information processed. A more compact model with fewer parameters not only avoids unnecessary computational and overfitting burdens but also achieves better scalability, embodying the “less is more” philosophy.
IV-F Larger Angular Resolution
IV-F1 Qualitative Comparison
The standard protocol of BasicLFSR [14] only engages the central $5\times 5$ SAIs. To explore the scalability of SkimLFSR to larger angular resolutions, we conduct an experiment on the $7\times 7$ SAIs to compare SkimLFSR with M2MT-Net [21] and LF-DET [9]. To handle the larger angular resolution, three DSA branches are employed in the Skim Transformers to handle the three corner SAI sets. The results are presented in the first group of Table IV.
SkimLFSR maintains its superiority over the other methods in this task as in the $5\times 5$ SAIs, achieving an average PSNR gain of 0.18 dB over the second-best method M2MT-Net. Particularly, on the EPFL and STFgantry datasets, SkimLFSR achieves a significant PSNR gain of 0.26 dB and 0.28 dB, respectively. This demonstrates that SkimLFSR is capable of handling different disparity ranges in a larger angular resolution.
IV-F2 Skimmed SAI Sets
As the angular resolution increases, the number of possible skimmed SAI combinations grows accordingly. As in Section IV-E3, we conduct an ablation study to evaluate the impact of different skimmed SAI sets. Using the three-branch SkimLFSR as the baseline, we construct three two-branch variants, (a), (b), and (c), each formed by removing one branch and thus retaining one of the three possible pairs of corner SAI sets. The results are presented in the second group of Table IV.
Overall, a consistent trend of average performance degradation is observed across the two-branch variants, ranging from 0.08 dB to 0.11 dB, relative to the three-branch baseline. While the overall degradation is limited, larger drops occur on specific datasets. Variants (a) and (c) fall behind by 0.15 dB on STFgantry, Variant (b) lags by 0.17 dB on INRIA, and Variant (c) shows a 0.18 dB decrease on EPFL.
These observations suggest that SkimLFSR benefits from employing more DSA branches when operating at higher angular resolutions, as additional branches provide more coverage of disparity ranges and consequently yield more robust reconstruction performance.
IV-F3 Cross-angular-resolution Generalization
As discussed in Section III-D4, SkimLFSR is inherently angular-resolution-agnostic. In fact, many existing LFSR methods are also angular-resolution-agnostic.
We validate this property by directly inferring models trained on the $5\times 5$ LF data to the $7\times 7$ LF data, without any retraining or major network modification. LFT [29], EPIT [28], LF-DET [9], and LFMamba [51] are included in this study. DistgSSR [46], HLFSR [40] and M2MT-Net [21] are not applicable for the reasons mentioned in Section III-D4. At the $7\times 7$ angular resolution, the model has access to a larger angular space. Thus, the two-branch SkimLFSR variant has the freedom to choose among three corner SAI sets to attend to. We therefore construct three variants, (d), (e), and (f), mirroring the same corner SAI sets as Variants (a), (b), and (c) in the previous study. Results are shown in the third group of Table IV.
It is interesting to observe that EPIT, which underperforms many existing methods on the $5\times 5$ SAIs task in Table II, surpasses them on the $7\times 7$ SAIs task. In particular, EPIT outperforms the Mamba-based method LFMamba by an average margin of 0.32 dB, whereas it lags behind by 0.21 dB on the $5\times 5$ setting.
For SkimLFSR, an even more surprising result is observed. Despite never being trained on $7\times 7$ LF data, the $5\times 5$ SkimLFSR variant achieves performance that is only up to 0.06 dB lower than the two-branch SkimLFSR variants in the second group that are trained on the $7\times 7$ LF data, and the improvement of training on $7\times 7$ mostly comes from the STFgantry dataset, bringing the PSNR score from 32.66-32.71 dB to 32.81 to 32.85 dB. It is also noteworthy that the generalized $5\times 5$ SkimLFSR model still outperforms both LF-DET and M2MT-Net in the first group, even though those methods in the first group are trained on $7\times 7$ LF data as well.
Motivated by these counter-intuitive observations, we hypothesize that a method’s angular generalizability is not necessarily correlated with its performance at a specific angular resolution. We attribute SkimLFSR’s strong performance in this study to the ability of its DSA branches to capture essential disparity information that is angular-resolution-agnostic, rather than cues that are tightly coupled to a particular angular configuration.
V Conclusion and Future Work
@@
In this paper, we presented SkimLFSR, a novel and efficient LFSR network that overcomes the challenge of disparity entanglement in LF image processing. The experimental results validated SkimLFSR’s “less is more” philosophy, demonstrating that SkimLFSR outperformed the state-of-the-art methods significantly by utilizing less but more pertinent information with reduced memory and computational complexity. Several analyses were conducted to reveal the underlying disparity-aware mechanisms and cross-angular-resolution generalizability of SkimLFSR.
Despite its success, our work has certain potential limitations that are worth further research. First, the skimmed SAI subsets of the Skim Transformers are manually pre-defined. Although this is not a complex problem and is only needed once for each model, it essentially introduces a prior into the network, which could potentially limit the generalization ability when it comes to a new angular resolution or an irregular angular subspace. One potential solution is to make the process of SAI selection heuristic or learned in network optimization autonomously. Second, most LF images in the current BasicLFSR [14] benchmark feature relatively simple objects and depth structures with limited intra-scene disparity variation. The light field community lacks systematic studies on how LFSR methods handle LF images with complex depth ranges within the same scene. Addressing this gap requires developing a dedicated dataset and evaluation protocol to systematically assess LFSR performance under complex intra-scene disparity variations in our future work. Third, while LF cameras are typically designed with regular micro-lens layouts, resulting in uniform grid-structured angular sampling, extending SkimLFSR to handle non-uniform angular spaces remains an open challenge. The reliance on regular network operations, such as standard convolutions that assume uniform spatial relationships, presents a potential obstacle to this extension. An additional consideration is whether current LFSR methods can effectively handle the variable disparities that arise from non-uniform angular sampling across different SAIs within the same LF image. Addressing these challenges may require the development of adaptive architectures capable of processing irregularly sampled angular data.
VI Acknowledgment
This work was supported in part by the Natural Science Foundation of China (No. 62577004) and the Open Project Program of State Key Laboratory of Virtual Reality Technology and Systems, Beihang University (VRLAB2025C04).